OpenStudy (anonymous):

121c5d7 (Under a radical) equals?

OpenStudy (anonymous):

11 c^5/2 d^7/2

OpenStudy (anonymous):

Is that formatting meaning c to the 5th over 2?

OpenStudy (anonymous):

\[ \sqrt{121c^5d^7} = 11c^{\frac{5}{2}}d^{\frac{7}{2}}\]

OpenStudy (anonymous):

Thank you:)